# In Praise of Stubbornness: The Case for Cognitive-Dissonance-Aware Knowledge Updates in LLMs

### v1

@@ A.2 Knowledge editing
Geva et al. (2020) were among the first to show that transformer Feed-Forward Network (FFN) layers act as unnormalized key-value stores encoding relational knowledge inside LLMs. This observation was later confirmed and complemented by others (Meng et al., 2022a; Dai et al., 2021) before being leveraged by subsequent work to master the editing of internal memories. Meng et al. (2022a) introduced ROME (Rank-One Model Editing), a method that uses causal tracing to empirically locate the layers essential to encoding a given association. They then modify these modules by applying small rank-one changes. To identify the relevant modules, they run the network multiple times, introducing corruptions to the input sequence to disturb the inference, and then restore individual states from the original non-corrupted pass. But this work an others worked only on single edits, and were often evaluated one edit at a time, starting each time from a fresh pre-trained model. The same authors later developed MEMIT, which follows the same causal tracing principle but with the goal of scaling up to 10,000 edits in bulk(Meng et al., 2022c). Similarly, Dai et al. (2021) leveraged the identification of knowledge neurons to perform “knowledge surgery” – editing factual knowledge within Transformers without the need for additional fine-tuning. Zhu et al. (2020) approached the knowledge modification task as a constrained optimization problem. Their work found that constrained layer-wise fine-tuning emerges as an effective method for modifying the knowledge that Transformers learn, suggesting a different pathway for knowledge editing inside LLMs. De Cao et al. (2021) proposed KnowledgeEditor, which achieved knowledge editing by training a hyper-network with constrained optimization to modify specific facts without fine-tuning or changing the overall stored knowledge. The method was demonstrated on smaller models like BERT for fact-checking and BART for question answering, achieving consistent changes in predictions across different formulations of queries.
@@
Li et al. (2023) empirically investigate the pitfalls of knowledge editing in LLMs, revealing two critical issues: logical inconsistencies between multiple edits (like contradictory relationship updates) and knowledge distortion (where edits irreversibly damage the model’s knowledge structure). Through carefully designed benchmarks ConflictEdit and RoundEdit, they demonstrate that current editing methods struggle with these challenges, particularly when handling reverse relationships or composite logical rules. While their work focuses on identifying limitations in maintaining logical consistency across edits, our paper takes a complementary cognitive-inspired perspective by addressing how models handle contradictions with their existing knowledge base. Their findings about knowledge distortion align with and reinforce our observations about the catastrophic nature of updates that modify existing knowledge.
@@
Similarly, Huang et al. (2024) empirically investigate causes of performance degradation during knowledge editing in LLMs. They show degradation correlates with editing target complexity and L1-norm growth in edited layers. Their proposed Dump for Sequence (D4S) method regulates layer norm growth using O(1) space complexity, enabling multiple effective updates while minimizing model degradation. Their work provides valuable insights into the mechanisms of model degradation during knowledge editing, but it does not specifically address the distinction between contradictory and non-contradictory updates, as we do in this paper.
@@
Tan et al. (2023) propose MALMEN, a scalable hypernetwork approach for editing Large Language Models by aggregating parameter shifts using a least-squares formulation. While previous editing methods like MEND (Mitchell et al., 2022) could handle only a few facts simultaneously, MALMEN can efficiently edit thousands of facts while maintaining comparable performance. Their key innovation lies in separating the computation between the hypernetwork and LM, enabling arbitrary batch sizes and reducing memory requirements. Their empirical results show that MALMEN can edit hundreds of times more facts than MEND while maintaining similar performance levels, though they note that the method still struggles with generalizing to rephrasing not seen during training. Like other editing approaches, MALMEN focuses on the mechanics of (by design conflicting) updates.
@@
Unlike all the work above, our goal in this work is not to edit knowledge, but to understand the fundamental mechanisms and phenomena that govern how LLMs integrate new information with existing knowledge. By taking a cognitive-inspired approach focused on dissonance awareness and adaptive plasticity, we reveal critical insights about the nature of knowledge representation and updating in these models.
@@ Appendix B Extraction of historical activations and gradients
We here detail our procedure for the extraction of activations and gradients. Source code is also available at https://github.com/bendiogene/ConflictAwareLLM/ for ultimate level of details and reproducibility purposes.
@@ C.4 Location of stubborn neurons
We also report the distribution of stubborn neurons across the transformer blocks in GPT-2 XL. Figures 8(a) and 8(b) show histograms of the number of stubborn neurons identified in each block for thresholds of 8,000 and 2,000 neurons, respectively.
@@ D.1 Similarities with Lottery ticket
To assess the hypothesis that certain subnetworks within the language model are more conducive to integrating new information—a notion earlier named the lottery ticket hypothesis (Frankle & Carbin, 2018)—we designed an experiment to confirm this effect.
@@ D.3 Comprehensive Analysis of GPT2-XL non-dissonant Updates
Our results reveal distinct scaling behaviors compared to GPT-2 small. With the optimal learning rate for GPT-2 XL (Figures 10(a)10(b)), we observe improved new knowledge acquisition while still preserving old knowledge. This means that although our carefully picked learning rate allows for efficient learning with full finetuning, learning with fewer neurons (as per our targetted strategies) seems harder than it was for GPT-2 small.
Increasing the learning rate by 10x (Figures 10(c)10(d)) or allocating 10x more neurons (Figures 10(e)10(f)) confirms that GPT-2 XL requires either higher learning rates or more extensive parameter updates compared to GPT-2 small to achieve effective learning with our targeted strategies.
@@
Similarly, extended training duration (50 epochs, Figures 10(g)10(h)) allows the model to better integrate new knowledge while preserving old information, indicating that longer training can also help overcome the limitations of sparse updates in larger models. Figure 11 summarizes these trade-offs across all configurations, highlighting how different hyperparameter choices affect the balance between preserving old knowledge and acquiring new information.
@@ E.1 Impact of number of conflicting facts
Our findings show that as the number of conflicting facts increases, the impact on old knowledge retention becomes more pronounced, with all strategies experiencing significant degradation. The ability to learn new conflicting knowledge improves slightly with more facts, but overall performance remains suboptimal. The plastic and random neuron strategies tend to preserve old knowledge when editing a small number of facts (e.g., 10 facts), but their effectiveness diminishes as more conflicting information is introduced. Interestingly, the opposite effect is observed for new knowledge, where adding more facts seems to make it easier to learn new knowledge, for all strategies.
@@
Figure 12: Dissonant updates with GPT2-small - impact of the number of conflicting facts. Each row represents a distinct metric: accuracy on the Generalization side dataset (paraphrased versions of the new facts), accuracy on New Knowledge, and Accuracy on Old Knowledge. Within each row, the subplots correspond to the number of conflicting facts introduced (10 Facts, 100 Facts, and 1000 Facts).
@@ E.2 Comparative performance of Editing methods
Our primary focus in this work is not on developing new model editing techniques. Most existing editing techniques focus on altering existing associations, and are hence by our definition dissonant by design. Our empirical findings in this work suggest another parallel path in which editing is abandoned in favor of non-dissonant variations where old knowledge is kept and contextualized
However, to have an idea on how existing editing methods perform compared to our targeted strategies, we leverage EasyEdit (Wang et al., 2023) to benchmark two state-of-the-art model editing methods, ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022c), under our same multi-fact experimental conditions .
Table 7 summarizes the performance of different strategies and editing methods. Some of our targeted update strategies obtain a higher harmonic mean compared to ROME and MEMIT. But the higher harmonic mean must not hide that the approaches are not directly comparable since they explore different regions of the pareto front, balancing new knowledge acquisition and old knowledge retention, as self-explained with colors and rankings in the table.
@@
MEMIT(Meng et al., 2022c) 0.962 (0.079)${}^{~{}1}$ 0.000 (${}^{~{}~{}}$ 0.000 (${}^{~{}~{}}$ 0.000 (${}^{~{}~{}}$
ROME(Meng et al., 2022a) 0.891 (${}^{~{}~{}}$ 0.240 (${}^{~{}~{}}$ 0.180 (${}^{~{}~{}}$ 0.236 (${}^{~{}~{}}$
@@
MEMIT(Meng et al., 2022c) 0.976 (0.008)${}^{~{}1}$ 0.004 (${}^{~{}~{}}$ 0.010 (${}^{~{}~{}}$ 0.003 (${}^{~{}~{}}$
ROME(Meng et al., 2022a) 0.431 (${}^{~{}~{}}$ 0.300 (${}^{~{}~{}}$ 0.150 (${}^{~{}~{}}$ 0.240 (${}^{~{}~{}}$
@@
MEMIT(Meng et al., 2022c) 0.605 (0.107)${}^{~{}1}$ 0.198 (${}^{~{}~{}}$ 0.100 (${}^{~{}~{}}$ 0.177 (${}^{~{}~{}}$
ROME(Meng et al., 2022a) 0.152 (${}^{~{}~{}}$ 0.160 (${}^{~{}~{}}$ 0.067 (${}^{~{}~{}}$ 0.106 (${}^{~{}~{}}$
@@ E.3 More detailed figures for specific numbers of neurons
The results show that changing the number of neurons allocated for updates does not necessarily improve or degrade performance in the dissonant update scenario. In all cases, the model struggles to retain old knowledge while learning new conflicting information. The candidate and specific neuron strategies are consistently and significantly better than state of the art solutions, offering a slight advantage. However, they are still unable to effectively mitigate the destructive effects of dissonant updates, further motivating the neeed for both (i) dissonance awareness and (ii) proper conflict resolution.
@@ E.4 Scaling to GPT2-XL
Old Knowledge New Knowledge Generalization
@@
(a) Old Knowledge
(b) New Knowledge
(c) Generalization
(d) Old Knowledge
(e) New Knowledge
(f) Generalization
(g) Old Knowledge
(h) New Knowledge
(i) Generalization
Figure 13: Dissonant updates with GPT2-XL: whether the model learns new knowledge or not, old knowledge is severely destroyed regardless of the strategy Experiments with 1000 facts using the best learning rate we found for Full Finetuning.
First, while GPT2-XL still requires more neurons than GPT-2 small to effectively learn new conflicting knowledge, as seen earlier, the key finding concerns old knowledge retention: regardless of model size or neuron allocation, we observe significant degradation of old, unrelated knowledge across all strategies.

### v2

@@ A.2 Knowledge editing
Geva et al. [2020] were among the first to show that transformer Feed-Forward Network (FFN) layers act as unnormalized key-value stores encoding relational knowledge inside LLMs. This observation was later confirmed and complemented by others [Meng et al., 2022a, Dai et al., 2021] before being leveraged by subsequent work to master the editing of internal memories. Meng et al. [2022a] introduced ROME (Rank-One Model Editing), a method that uses causal tracing to empirically locate the layers essential to encoding a given association. They then modify these modules by applying small rank-one changes. To identify the relevant modules, they run the network multiple times, introducing corruptions to the input sequence to disturb the inference, and then restore individual states from the original non-corrupted pass. But this work an others worked only on single edits, and were often evaluated one edit at a time, starting each time from a fresh pre-trained model. The same authors later developed MEMIT, which follows the same causal tracing principle but with the goal of scaling up to 10,000 edits in bulk[Meng et al., 2022c]. Similarly, Dai et al. [2021] leveraged the identification of knowledge neurons to perform “knowledge surgery" – editing factual knowledge within Transformers without the need for additional fine-tuning. Zhu et al. [2020] approached the knowledge modification task as a constrained optimization problem. Their work found that constrained layer-wise fine-tuning emerges as an effective method for modifying the knowledge that Transformers learn, suggesting a different pathway for knowledge editing inside LLMs. De Cao et al. [2021] proposed KnowledgeEditor, which achieved knowledge editing by training a hyper-network with constrained optimization to modify specific facts without fine-tuning or changing the overall stored knowledge. The method was demonstrated on smaller models like BERT for fact-checking and BART for question answering, achieving consistent changes in predictions across different formulations of queries.
@@
Li et al. [2023] empirically investigate the pitfalls of knowledge editing in LLMs, revealing two critical issues: logical inconsistencies between multiple edits (like contradictory relationship updates) and knowledge distortion (where edits irreversibly damage the model’s knowledge structure). Through carefully designed benchmarks ConflictEdit and RoundEdit, they demonstrate that current editing methods struggle with these challenges, particularly when handling reverse relationships or composite logical rules. While their work focuses on identifying limitations in maintaining logical consistency across edits, our paper takes a complementary cognitive-inspired perspective by addressing how models handle contradictions with their existing knowledge base.
@@
Similarly, Huang et al. [2024] empirically investigate causes of performance degradation during knowledge editing in LLMs. They show degradation correlates with editing target complexity and L1-norm growth in edited layers. Their proposed Dump for Sequence (D4S) method regulates layer norm growth using O(1) space complexity, enabling multiple effective updates while minimizing model degradation. Their work provides valuable insights into the mechanisms of model degradation during knowledge editing.
@@
Tan et al. [2023] propose MALMEN, a scalable hypernetwork approach for editing Large Language Models by aggregating parameter shifts using a least-squares formulation. While previous editing methods like MEND [Mitchell et al., 2022] could handle only a few facts simultaneously, MALMEN can efficiently edit thousands of facts while maintaining comparable performance. Their key innovation lies in separating the computation between the hypernetwork and LM, enabling arbitrary batch sizes and reducing memory requirements. Their empirical results show that MALMEN can edit hundreds of times more facts than MEND while maintaining similar performance levels, though they note that the method still struggles with generalizing to rephrasing not seen during training. Like other editing approaches, MALMEN focuses on the mechanics of (by design conflicting) updates.
@@
Unlike all the work above, our goal in this work is not to edit existing knowledge (which is conflicting by design according to our definition), but to understand the fundamental mechanisms and phenomena that govern how LLMs integrate new information with existing knowledge, contrasting contradictory and non-contradictory updates. By taking inspiration from humans’ cognitive-dissonance, we reveal critical insights about the nature of knowledge representation and updating in these models.
A.3 Related empirical and theoretical findings
To the best of our knowledge, we are the first to report the systematic catastrophic effect of contradictory updates on completely unrelated knowledge in LLMs (in addition to the benefits of differentiated plasticity in case of non-dissonant updates). A couple of contemporary findings are to an extent related to ours.
First, Hu et al. [2025] have recently shown that finetuning models to write insecure code leads to emergent misalignment across unrelated domains. Through the lens of our work, their findings may represent another manifestation of how contradictory updates propagate through neural networks. By training models to produce insecure code without disclosure (effectively contradicting their ethical alignment training) they observed widespread behavioral changes far beyond the coding domain, including anti-human viewpoints and deceptive behaviors. Similar to how our factual contradictions (e.g., changing nationality information) destroy unrelated factual knowledge, their ethical contradictions appear to disrupt the model’s broader behavioral alignment. This parallel suggests that the catastrophic interference we document may be a more fundamental property of neural networks than previously recognized, affecting not only factual knowledge but potentially also learned ethical constraints.
Second, not related to LLMs but worth mentioning, Hiratani [2024] have analytically uncovered an interesting asymmetry in continual learning using a simple model as a playground for tractability (a linear teacher-student model with latent structure). They demonstrated analytically that this model struggles catastrophically when familiar inputs must be mapped to entirely new outputs (high input similarity, low output similarity), while performing relatively well in the opposite scenario (low input similarity, high output similarity). They further validated some key predictions using a single-hidden-layer network on a permuted MNIST task, but acknowledged that deeper networks might enable different adaptations to feature and readout similarity, calling for future work to adapt to more complex architectures and scenarios. Our empirical findings with various LLM models and facts reveal a potentially related phenomenon, suggesting an asymmetry that extends beyond this simple model. In fact, when we introduce contradictory updates (e.g. changing "Danielle Darrieux is French" to "English"), we observe catastrophic corruption of unrelated knowledge - conceptually similar to the high-input/low-output similarity scenario. Conversely, non-contradictory updates cause minimal interference. This parallel is particularly noteworthy given the vast differences in architecture complexity, domain, and scale between their controlled experiments and our work with LLMs. More research is needed to formalize this connection and determine under which conditions this represents a universal property of neural network learning more generally.
Finally, in a concurrent work, Hu et al. [2025] mathematically analyzed knowledge editing methods that use linear associative memory (like ROME and MEMIT) which directly modify specific weights rather than using gradient descent. They proved that these methods will always inevitably suffer from interference due to knowledge superposition in the model’s parameter space. While their analysis is specific to these specialized editing techniques and primarily demonstrated with semantically related concepts (e.g., ’Vladimir Mayakovsky’ and ’Vladimir Bukovsky’ are highly superposed, meaning that editing one, will interfere with the other), our work reveals a parallel but distinct phenomenon in conventional gradient-based learning: contradictory updates cause catastrophic corruption of completely unrelated knowledge. This suggests that while the specific mechanisms differ between direct weight editing and gradient-based learning, both approaches encounter fundamental limitations when modifying existing knowledge. Our discovery of the stark asymmetry between contradictory and non-contradictory updates provides a complementary perspective to their superposition analysis, suggesting that the nature of the update itself is a critical factor in determining interference patterns across different training paradigms. A promising direction for future work would be to develop an equivalent “superposition” framework for gradient-based learning, potentially investigating whether contradictory updates create more disruptive patterns in the model’s representation space than non-contradictory ones.
@@ Appendix B Extraction of historical activations and gradients
We here detail our procedure for the extraction of activations and gradients. Source code is also available at https://figshare.com/s/81f7108d823b5e08e8ec for ultimate level of details and reproducibility purposes.
@@ C.4 Location of stubborn neurons
We also report the distribution of stubborn neurons across the transformer blocks in GPT-2 XL. Figures 7(a) and 7(b) show histograms of the number of stubborn neurons identified in each block for thresholds of 8,000 and 2,000 neurons, respectively.
@@ D.1 Similarities with Lottery ticket
To assess the hypothesis that certain subnetworks within the language model are more conducive to integrating new information—a notion earlier named the lottery ticket hypothesis [Frankle and Carbin, 2018]—we designed an experiment to confirm this effect.
@@ D.3 Comprehensive Analysis of GPT2-XL non-dissonant Updates
Our results reveal distinct scaling behaviors compared to GPT-2 small. With the optimal learning rate for GPT-2 XL (Figures 9(a)9(b)), we observe improved new knowledge acquisition while still preserving old knowledge. This means that although our carefully picked learning rate allows for efficient learning with full finetuning, learning with fewer neurons (as per our targetted strategies) seems harder than it was for GPT-2 small.
Increasing the learning rate by 10x (Figures 9(c)9(d)) or allocating 10x more neurons (Figures 9(e)9(f)) confirms that GPT-2 XL requires either higher learning rates or more extensive parameter updates compared to GPT-2 small to achieve effective learning with our targeted strategies.
@@
Similarly, extended training duration (50 epochs, Figures 9(g)9(h)) allows the model to better integrate new knowledge while preserving old information, indicating that longer training can also help overcome the limitations of sparse updates in larger models. Figure 10 summarizes these trade-offs across all configurations, highlighting how different hyperparameter choices affect the balance between preserving old knowledge and acquiring new information.
@@ E.1 Impact of number of conflicting facts
Our findings show that as the number of conflicting facts increases, the impact on old unrelated knowledge retention becomes more pronounced, with all strategies experiencing significant degradation. The ability to learn new conflicting knowledge improves slightly with more facts, but overall performance remains suboptimal. The plastic and random neuron strategies tend to preserve old knowledge when editing a small number of facts (e.g., 10 facts), but their effectiveness diminishes as more conflicting information is introduced. Interestingly, the opposite effect is observed for new knowledge, where adding more facts seems to make it easier to learn new knowledge, for all strategies.
@@
Figure 11: Dissonant updates with GPT2-small - impact of the number of conflicting facts. Each row represents a distinct metric: accuracy on the Generalization side dataset (paraphrased versions of the new facts), accuracy on New Knowledge, and Accuracy on old unrelated knowledge. Within each row, the subplots correspond to the number of conflicting facts introduced (10 Facts, 100 Facts, and 1000 Facts).
@@ E.2 Comparative performance of editing methods
Our primary focus in this work is not on developing new model editing techniques. Most existing editing techniques focus on altering existing associations, and are hence by our definition dissonant by design. Our empirical findings in this work suggest another parallel path in which editing is abandoned in favor of non-dissonant variations where old knowledge is kept and contextualized.
However, to have an idea on how existing editing methods perform compared to our targeted strategies, we leverage EasyEdit [Wang et al., 2023] to benchmark two state-of-the-art model editing methods, ROME [Meng et al., 2022a] and MEMIT [Meng et al., 2022c], under our same multi-fact experimental conditions .
Table E.2 summarizes the performance of different strategies and editing methods. Some of our targeted update strategies obtain a higher harmonic mean compared to ROME and MEMIT. But the higher harmonic mean must not hide that the approaches are not directly comparable since they explore different regions of the pareto front, balancing new knowledge acquisition and old knowledge retention, as self-explained with colors and rankings in the table.
@@
MEMIT[Meng et al., 2022c] 0.962 (0.079)${}^{~{}1}$ 0.000 (${}^{~{}~{}}$ 0.000 (${}^{~{}~{}}$ 0.000 (${}^{~{}~{}}$
ROME[Meng et al., 2022a] 0.891 (${}^{~{}~{}}$ 0.240 (${}^{~{}~{}}$ 0.180 (${}^{~{}~{}}$ 0.236 (${}^{~{}~{}}$
@@
MEMIT[Meng et al., 2022c] 0.976 (0.008)${}^{~{}1}$ 0.004 (${}^{~{}~{}}$ 0.010 (${}^{~{}~{}}$ 0.003 (${}^{~{}~{}}$
ROME[Meng et al., 2022a] 0.431 (${}^{~{}~{}}$ 0.300 (${}^{~{}~{}}$ 0.150 (${}^{~{}~{}}$ 0.240 (${}^{~{}~{}}$
@@
MEMIT[Meng et al., 2022c] 0.605 (0.107)${}^{~{}1}$ 0.198 (${}^{~{}~{}}$ 0.100 (${}^{~{}~{}}$ 0.177 (${}^{~{}~{}}$
ROME[Meng et al., 2022a] 0.152 (${}^{~{}~{}}$ 0.160 (${}^{~{}~{}}$ 0.067 (${}^{~{}~{}}$ 0.106 (${}^{~{}~{}}$
@@ E.3 More detailed figures for specific numbers of neurons
The results show that changing the number of neurons allocated for updates does not necessarily improve or degrade performance in the dissonant update scenario. In all cases, the model struggles to retain old unrelated knowledge while learning new conflicting information. The candidate and specific neuron strategies are consistently and significantly better than state of the art solutions, offering a slight advantage. However, they are still unable to effectively mitigate the destructive effects of dissonant updates, further motivating the neeed for both (i) dissonance awareness and (ii) proper conflict resolution.
@@ E.4 Scaling to GPT2-XL
old unrelated knowledge New Knowledge Generalization
@@
(j) old unrelated knowledge
(k) New Knowledge
(l) Generalization
(m) old unrelated knowledge
(n) New Knowledge
(o) Generalization
(p) old unrelated knowledge
(q) New Knowledge
(r) Generalization
Figure 12: Dissonant updates with GPT2-XL: whether the model learns new knowledge or not, old unrelated knowledge is severely destroyed regardless of the strategy Experiments with 1000 facts using the best learning rate we found for Full Finetuning.
First, while GPT2-XL still requires more neurons than GPT-2 small to effectively learn new conflicting knowledge, as seen earlier, the key finding concerns unrelated knowledge retention: regardless of model size or neuron allocation, we observe significant degradation of old, unrelated knowledge across all strategies.
@@
Appendix F Limitations
F.1 Scaling dissonance beyond simple counterfacts
Our controlled experiments using COUNTERFACT’s simple factual statements (e.g., “Paris is the capital of France”) may not capture the full complexity of knowledge conflicts in real-world deployments. More complex forms of knowledge—procedural skills, reasoning patterns, coding tricks, ethical guidelines, or nuanced conceptual understanding—might exhibit different types of dissonance and hence interference patterns than the factual relations we studied. In this work, we showed that contradictory updates (that change known factual associations) corrupt unrelated knowledge
F.2 Contradiction Detection Generalizability
Also related, our 95%+ accuracy (98% when using output probabilities) in detecting contradictions was achieved on balanced datasets with clear fact/counterfact pairs. Real-world contradictions often involve much longer text, subtler conflicts, implicit assumptions, or context-dependent truths that may be harder to detect. The generalization of our detection methods to such nuanced scenarios remains untested. Future work should design dedicated datasets and benchmarks for conflict detection.
F.3 Scale Limitations
We showed the success of selective plasticity in case of non-dissonant updates. However, the relatively small scale of our tracked knowledge (2,000-3,000 facts) compared to the full knowledge capacity of larger models limits our ability to observe interference effects, a limit we starting already from GPT-2-XL (our 2000 facts represent an even tinier portion of the models knowledge).
For dissonant updates, while we observed the same consistent catastrophic patterns up to 6B parameters, it is unclear how these effects manifest in even larger models like today’s commercial LLMs.
Appendix G Impact Statement
These findings motivate our development of dissonance detection capabilities as a crucial safety mechanism for deployed AI systems. More broadly, they suggest we may need to fundamentally rethink AI architectures to develop systems that, like humans, maintain and contextualize potentially conflicting information rather than attempting to overwrite it. This might require moving away from current approaches that try to "edit" neural networks toward architectures that can accumulate and contextualize knowledge while maintaining multiple, temporally-organized versions of truth.